# Adapting In-Domain Few-Shot Segmentation to New Domains without Retraining

### v2

@@ Abstract
Cross-domain few-shot segmentation (CD-FSS) aims to segment objects of novel classes in new domains, which is often challenging due to the diverse characteristics of target domains and the limited availability of support data. Most CD-FSS methods redesign and retrain in-domain FSS models using various domain-generalization techniques, which are effective but costly to train. To address these issues, we propose adapting informative model structures of the well-trained FSS model for target domains by learning domain characteristics from few-shot labeled support samples during inference, thereby eliminating the need for retraining. Specifically, we first adaptively identify domain-specific model structures by measuring parameter importance using a novel structure Fisher score in a data-dependent manner. Then, we progressively train the selected informative model structures with hierarchically constructed training samples, progressing from fewer to more support shots. The resulting Informative Structure Adaptation (ISA) method effectively addresses domain shifts and equips existing well-trained in-domain FSS models with flexible adaptation capabilities for new domains, eliminating the need to redesign or retrain CD-FSS models on base data. Extensive experiments validate the effectiveness of our method, demonstrating superior performance across multiple CD-FSS benchmarks.
22footnotetext: Corresponding author: liunian228@gmail.com
@@ 1 Introduction
Few-shot semantic segmentation (FSS) aims to segment novel classes using a limited number of support samples. It typically trains a conventional support-query matching network to transfer class-agnostic patterns from extensive base data to novel classes. Existing FSS methods [11, 37, 35, 64] have made significant progress on in-domain class generalization due to various well-designed model architectures, as well as matching and training techniques.
@@
Figure 1: Cross-domain few-shot segmentation is usually challenging due to ➀ diverse characteristics of target domains (e.g., requiring high- or low-level parsing) and the ➁ limited support data (e.g., 1 to 5 shots). Our Informative Structure Adaptation (ISA) can adaptively identify and efficiently train domain-specific informative structures for the target domains during inference. This process is applied to few-shot annotated support data during inference and can be directly integrated with various well-trained FSS methods [11, 54, 66, 67, 45] without retraining, leading to substantial improvements in cross-domain few-shot segmentation.
@@
Existing CD-FSS methods [25, 45, 57, 19] typically train models using abundant base data from the source domain with various domain-generalization techniques to address the domain shifts problem. However, retraining on base data is costly and fails to leverage powerful in-domain FSS models. Besides, the issue arises because CD-FSS models are trained on limited domain data with frozen parameters, while the potential target domains can be diverse and arbitrary. Therefore, it is necessary to adapt the well-trained in-domain FSS models to target domains during inference.
@@
Test-time training (TTT) [52, 49] effectively adapts models to target domains by learning from test data. Although promising, most existing TTT methods adjust the same manually-selected trainable structures across different domains, disregarding the distinct domain gaps and characteristics of the target domains. However, different domains, or even individual test images, exhibit distinct properties related to various model structures, such as specific model layers. For example, the satellite images in DeepGlobe [7] dataset primarily rely on low-level texture analysis for parsing structured and detailed remote sensing areas. In contrast, Chest X-ray images [2] typically require middle- and high-level semantic understanding to distinguish pneumonia-affected areas from normal lung regions in medical image analysis. Thus, dynamically selecting trainable model structures is crucial for adapting to various target domains with distinct characteristics.
@@
To address these challenges, we propose a novel method, Informative Structure Adaptation (ISA), specifically designed for cross-domain few-shot segmentation. We explore methods for identifying and adapting domain-specific informative structures during inference by learning domain characteristics from few-shot labeled support samples. Given the varying reliance on model structures across different domains [60, 29], the Informative Structure Identification (ISI) module identifies domain-sensitive model structures by measuring parameter importance in a data-dependent manner. First, we compute empirical Fisher information to reduce computational overhead, inspired by the parameter importance metric used in continual learning [24]. Building on this, we propose a novel structure Fisher score to guide the identification of informative model structures. This strategy optimizes model adaptation for varying domain characteristics and mitigates the risk of overfitting in few-shot scenarios.
@@
Our approach, as shown in Figure 1, fundamentally differs from conventional few-shot segmentation and test-time training methods. We leverage few-shot annotated support data to adaptively identify and progressively adapt informative model structures during test-time training for CD-FSS. In contrast, most prevailing test-time training methods [52, 46, 55, 47] manually define trainable model structures and directly train models on accessible test data, which are proven suboptimal by our empirical analysis. Although PATNet [25] introduces a test-time training method for CD-FSS, it is specifically designed to train fixed anchor layers and requires pre-training on the source domain. Therefore, applying PATNet to other few-shot semantic segmentation methods is non-trivial. In contrast, our method is model-agnostic, requires no additional learnable parameters, and can easily equip existing few-shot segmentation models with flexible adaptation capabilities for new domains, eliminating the need to redesign or retrain CD-FSS models on base data. In summary, our key contributions include:
  • We propose, for the first time, a novel framework Informative Structure Adaptation (ISA) for CD-FSS that generalizes well-trained in-domain few-shot segmentation models to new domains without retraining.
@@ 2 Related Works
Few-Shot Semantic Segmentation. Few-shot semantic segmentation, pioneered by Shaban et al. [41], aims to predict dense masks for objects of novel classes using only a limited number of labeled support images. The mainstream prototype-based methods [8, 26, 54] perform segmentation by measuring the similarity between the query features and representative support prototypes incorporating various improvements [43, 31, 62, 70]. The affinity-based methods [35, 65, 39, 36, 50] establish detailed dense correspondence between query and support features through feature concatenation and leverage a learnable CNN or transformer module for segmentation prediction. Recently, foundation models like SAM [23] present a novel opportunity for few-shot segmentation [34, 67], due to their remarkable transfer capability on tasks and data distributions beyond the training scope. However, these methods do not consider the domain shifts problem, leading to poor generalization performance when encountering new domains during testing.
@@
Cross-Domain Few-Shot Semantic Segmentation. Cross-domain few-shot semantic segmentation has recently received increasing attention. PATNet [25] introduces a feature transformation layer that seamlessly maps query and support features across diverse domains into a unified feature space, effectively tackling the intra-domain knowledge preservation issue in CD-FSS. RD [57] utilizes a memory bank to reinstill meta-knowledge from the source domain, thereby improving generalization performance in the target domain. Subsequently, DARNet [10] and RestNet [19] approach the problem from distinct perspectives, focusing on bridging domain gaps through dynamic adaptation refinement and knowledge transfer, respectively. Inspired by these pioneering efforts, PMNET [3] presents a comprehensive solution capable of addressing both in-domain and cross-domain FSS tasks concurrently by capturing pixel relations within each support-query pair. Unlike previous CD-FSS approaches, our method effectively addresses domain shifts in CD-FSS during inference by adaptively identifying and gradually adapting informative model structures on few-shot annotated support samples, and is capable of seamlessly adapting current few-shot segmentation methods to address domain shifts problem.
@@
Test-Time Training. Normally, once a well-trained model is deployed, it remains static without further alterations. In contrast, test-time training (TTT) [48] adapts models to the deployment scenario by leveraging unlabeled data available at test time. The mainstream self-supervised learning-based methods [33, 53, 30, 14, 13] leverage the available unlabeled test data to facilitate model adaptation to the target domain using self-supervised learning techniques. The feature alignment-based methods [46, 22, 56] attempt to rectify the feature representations for the target domain. Some works attempt to apply TTT to address the semantic segmentation problem. For instance, MM-TTA [42] utilizes multiple modalities to provide reciprocal TTT self-supervision for 3D semantic segmentation. Similarly, CD-TTA [44] explores domain-specific TTT for urban scene segmentation using an online clustering algorithm. OCL [68] proposes an output contrastive loss to stabilize the TTT adaptation process for extreme class imbalance and complex decision spaces in semantic segmentation. These methods typically adjust fixed hand-selected trainable structures on one single test image for different domains. In contrast, our method dynamically adapt domain-specific informative structures by learning domain characteristics from few-shot labeled support samples.
@@ 3 Method
Cross-Domain Few-Shot Segmentation (CD-FSS) aims to transfer knowledge learned from the source domain to new categories in unseen target domains using minimal annotated support samples. The model is typically trained on the source domain and then evaluated on target domains, ensuring no label space overlap between the source and target domains.
@@ 3.1 Baseline Methods
Few-Shot Segmentation Model. Mainstream few-shot semantic segmentation model can be formulated as follows: The input support and query images $\{I_{s},I_{q}\}$ are processed by a weight-shared backbone to extract image features $\{\mathcal{F}_{s},\mathcal{F}_{q}\}$:
@@
where $f$ denotes the image encoder with parameters $\theta$. Then, the support features $\mathcal{F}_{s}$ and groundtruth masks $\mathcal{M}_{s}$ are fed into the masked average pooling layer (MAP) to generate support prototypes $\mathcal{P}_{s}$. The final prediction is made by measuring the cosine similarity between $\mathcal{P}_{s}$ and $\mathcal{F}_{q}$.
@@
3:for $n$ from $1$ to $K-1$ do $\triangleright$ PSA with increasing support shots. See Eqn. 7
4:     Extract features $\mathcal{F}$ for all samples using the updated model $\mathit{f}$triangleright See Eqn. 1
@@
9:     Compute average loss $\mathcal{L}_{\texttt{PSA},n}$ for support shots $n$triangleright See Eqn. 5
@@
11:     Update model $\mathit{f}$: $\theta_{\texttt{tr},n-1}^{*}$rightarrow $\theta_{\texttt{tr},n}^{*}$triangleright See Eqn. 6
@@
[…]
@@ 3.2 Informative Structure Identification
Empirical Fisher Information. However, directly computing the Fisher Information Matrix for the pre-trained CD-FSS model involves significant computational overhead due to the $|\theta|\times|\theta|$ scale. Thus, we simplify FIM for CD-FSS, inspired by the parameter importance metric used in continual learning [24]. Specifically, we concentrate on the support samples $K$ in the target domain and utilize the diagonal elements of the “Empirical Fisher” to evaluate the importance of the pre-trained model parameters for cross-domain tasks. Specifically, for the $l$-th convolutional layer, we derive the empirical Fisher information of its $u$-th trainable parameters as $F_{\theta_{l,u}}=\frac{1}{|K|}\sum_{j=1}^{|K|}\left(\frac{\partial\log p_{\theta}(y_{j}|x_{j})}{\partial\theta_{l,u}}\right)^{2}$. Correspondingly, a relatively large value of $F_{\theta_{l,u}}$ indicates that the parameter $\theta_{l,u}$ is crucial for the cross-domain task.
@@
Model layers with higher structure Fisher scores are typically more important for model training [32] because of their greater contribution to the optimization process. Updating only the informative model structures preserves the model’s ability to fit few-shot data and regularizes training to mitigate the risk of overfitting. Therefore, we select the model layer with the highest structure Fisher score and update its parameters $\theta_{\texttt{tr}}$ for model structure adaptation during inference, while freezing all other parameters to minimize the risk of overfitting in few-shot scenarios:
[…]
@@ 3.3 Progressive Structure Adaptation
Table 1: Quantitative comparison results on the CD-FSS benchmark. The models are trained on the Pascal VOC source domain dataset and evaluated on four datasets exhibiting distinct domain shifts. The best results are highlighted with bold. The $\dagger$ means our reproduced results. The $\ddagger$ means using the ViT-base backbone. All the in-domain FSS methods are directly applied to CD-FSS without retraining (✗), thus suffering from the domain shifts problem. All other CD-FSS methods requires retraining (✓) their models on the source domain to enhance domain generalization. Our method can be directly integrated with the well-trained FSS method (e.g., SSP) without retraining (✗).
@@
PGNet [63] ICCV 2019 10.7 12.4 21.9 21.3 34.0 23.0 62.4 62.7 32.2 31.1
PANet [54] ICCV 2019 36.6 45.4 25.3 34.0 57.8 69.3 69.2 71.7 47.2 55.1
CaNet [64] CVPR 2019 22.3 23.1 25.2 28.2 28.4 28.6 70.7 72.0 36.6 38.0
RPMMs [59] ECCV 2020 13.0 13.5 18.0 20.0 30.1 30.8 65.1 67.1 31.6 32.9
PFENet [50] TPAMI 2020 16.9 18.0 23.5 23.8 27.2 27.6 70.9 70.5 34.6 35.0
RePRI [1] CVPR 2021 25.0 27.4 23.3 26.2 65.1 65.5 71.0 74.2 46.1 48.3
HSNet [36] ICCV 2021 29.7 35.1 31.2 35.1 51.9 54.4 77.5 81.0 47.6 51.4
SSP [11] ECCV 2022 42.3 50.4 33.0 47.0 74.9 75.5 77.1 79.1 56.8 63.0
PATNet [25] ECCV 2022 37.9 43.0 41.2 53.6 66.6 70.2 78.6 81.2 56.1 62.0
PMNet [3] WACV 2024 37.1 41.6 51.2 54.5 70.4 74.0 84.6 86.3 60.8 64.1
ABCDFSS [18] CVPR 2024 42.6 49.0 45.7 53.3 79.8 81.4 74.6 76.2 60.7 65.0
APSeg [17] CVPR 2024 35.9 40.0 45.4 54.0 84.1 84.5 79.7 81.9 61.3 65.1
DR-Adapter [45] CVPR 2024 41.3 50.1 40.8 48.9 82.4 82.3 79.1 80.4 60.9 65.4
@@
Hierarchical Training Sample Construction. We begin by enhancing the utilization of few-shot support data, cyclically designating each support image as a pseudo query image to generate multiple support-query training pairs $\{(S\setminus S^{i}_{q},S^{i}_{q})\}_{i=1}^{K}$. To conserve computational resources, we first extract features from all support images, then compute the losses for each constructed training pair based on these features, and finally perform back-propagation to update model parameters using the averaged loss $\frac{1}{K}\sum_{i=1}^{K}\mathcal{L}_{\texttt{T}}^{i}$ across all training pairs. Next, we vary the number of support shots from $1$ to $K-1$ to construct hierarchical training pairs. Specifically, for each support shot number $n$ ($n\leq K-1$), we construct training pairs containing $n$ support samples. Thus, we progressively increase the number of support samples to construct training pairs, optimizing the utilization of few-shot support data and enabling the model to gradually handle domain shifts during inference.
Progressive Structure Adaptation. To better address domain shifts in cross-domain tasks during test-time training, we propose a Progressive Structure Adaptation (PSA) module that trains the model using HTC-constructed hierarchical training pairs. The PSA method progressively trains the model with an increasing number of support shots from $1$ to $K-1$, gradually reducing the domain gap. Specifically, for the support shot number $n$, the training loss is:
@@
where $\theta_{n-1}^{*}$ and $\theta_{n}^{*}$ denote the updated model parameters trained with HTC-constructed training samples with $n-1$ and $n$ support shots, respectively. Specifically, when $n=1$, the $\theta_{n-1}^{*}$ represents the original parameters $\theta$ of the FSS model.
Consequently, we progressively update the model parameters by gradually increasing the number of support shots from $1$ to $K-1$, as follows: $\theta_{1}^{*}\rightarrow\theta_{2}^{*}\rightarrow\dots\rightarrow\theta_{n}^{*}\rightarrow\dots\rightarrow\theta_{K-1}^{*}.$ This approach optimizes the use of limited support data to progressively mitigate domain shifts during inference on the support set. Hence, we term it as Progressive Structure Adaptation (PSA) module.

### v3

# Adapting In-Domain Few-Shot Segmentation to New Domains without Source Domain Retraining

@@ Abstract
Cross-domain few-shot segmentation (CD-FSS) aims to segment objects of novel classes in new domains, which is often challenging due to the diverse characteristics of target domains and the limited availability of support data. Most CD-FSS methods redesign and retrain in-domain FSS models using abundant base data from the source domain, which are effective but costly to train. To address these issues, we propose adapting informative model structures of the well-trained FSS model for target domains by learning domain characteristics from few-shot labeled support samples during inference, thereby eliminating the need for source domain retraining. Specifically, we first adaptively identify domain-specific model structures by measuring parameter importance using a novel structure Fisher score in a data-dependent manner. Then, we progressively train the selected informative model structures with hierarchically constructed training samples, progressing from fewer to more support shots. The resulting Informative Structure Adaptation (ISA) method effectively addresses domain shifts and equips existing well-trained in-domain FSS models with flexible adaptation capabilities for new domains, eliminating the need to redesign or retrain CD-FSS models on base data. Extensive experiments validate the effectiveness of our method, demonstrating superior performance across multiple CD-FSS benchmarks. Codes are at https://github.com/fanq15/ISA.
@@ 1 Introduction
Few-shot semantic segmentation (FSS) aims to segment novel classes using a limited number of support samples. It typically trains a conventional support-query matching network to transfer class-agnostic patterns from extensive base data to novel classes. Existing FSS methods [fan2022self, nguyen2019feature, lu2021simpler, zhang2019canet] have made significant progress on in-domain class generalization due to various well-designed model architectures, as well as matching and training techniques.
@@
Figure 1: Cross-domain few-shot segmentation is usually challenging due to ➀ diverse characteristics of target domains (e.g., requiring high- or low-level parsing) and the ➁ limited support data (e.g., 1 to 5 shots). Our Informative Structure Adaptation (ISA) can adaptively identify and efficiently train domain-specific informative structures for the target domains during inference. This process is applied to few-shot annotated support data during inference and can be directly integrated with various well-trained FSS methods [fan2022self, wang2019panet, zhang2022feature, zhang2023personalize, su2024domain] without source domain retraining, leading to substantial improvements in CD-FSS.
@@
Existing CD-FSS methods [lei2022cross, su2024domain, wang2022remember, huang2023restnet] typically train models using abundant base data from the source domain with various domain-generalization techniques to address the domain shifts problem. However, retraining on source domain is costly and fails to leverage powerful in-domain FSS models. Besides, the issue arises because CD-FSS models are trained on limited domain data with frozen parameters, while the potential target domains can be diverse and arbitrary. Therefore, it is necessary to adapt the well-trained in-domain FSS models to target domains during inference.
@@
Test-time training (TTT) [wang2020tent, sun2020test] effectively adapts models to target domains by learning from test data. Although promising, most existing TTT methods adjust the same manually-selected trainable structures across different domains, disregarding the distinct domain gaps and characteristics of the target domains. However, different domains, or even individual test images, exhibit distinct properties related to various model structures, such as specific model layers. For example, the satellite images in DeepGlobe [demir2018deepglobe] dataset primarily rely on low-level texture analysis for parsing structured and detailed remote sensing areas. In contrast, Chest X-ray images [candemir2013lung] typically require middle- and high-level semantic understanding to distinguish pneumonia-affected areas from normal lung regions in medical image analysis. Thus, dynamically selecting trainable model structures is crucial for adapting to various target domains with distinct characteristics.
@@
To address these challenges, we propose a novel method, Informative Structure Adaptation (ISA), specifically designed for cross-domain few-shot segmentation. We explore methods for identifying and adapting domain-specific informative structures during inference by learning domain characteristics from few-shot labeled support samples. Given the varying reliance on model structures across different domains [yosinski2014transferable, liang2019distant], the Informative Structure Identification (ISI) module identifies domain-sensitive model structures by measuring parameter importance in a data-dependent manner. First, we compute empirical Fisher information to reduce computational overhead, inspired by the parameter importance metric used in continual learning [kirkpatrick2017overcoming]. Building on this, we propose a novel structure Fisher score to guide the identification of informative model structures. This strategy optimizes model adaptation for varying domain characteristics and mitigates the risk of overfitting in few-shot scenarios.
@@
Our approach, as shown in Figure 1, fundamentally differs from conventional few-shot segmentation and test-time training methods. We leverage few-shot annotated support data to adaptively identify and progressively adapt informative model structures during test-time training for CD-FSS. In contrast, most prevailing test-time training methods [wang2020tent, su2022revisiting, wang2022continual, su2024towards] manually define trainable model structures and directly train models on accessible test data, which are proven suboptimal by our empirical analysis. Although PATNet [lei2022cross] introduces a test-time training method for CD-FSS, it is specifically designed to train fixed anchor layers and requires pre-training on the source domain. Therefore, applying PATNet to other few-shot semantic segmentation methods is non-trivial. In contrast, our method is model-agnostic, requires no additional learnable parameters, and can easily equip existing few-shot segmentation models with flexible adaptation capabilities for new domains, eliminating the need to redesign or retrain CD-FSS models on base data. In summary, our key contributions include:
  • We propose, for the first time, a novel framework Informative Structure Adaptation (ISA) for CD-FSS that generalizes well-trained in-domain few-shot segmentation models to new domains without source domain retraining.
@@ 2 Related Works
Few-Shot Semantic Segmentation. Few-shot semantic segmentation, pioneered by Shaban et al. [shaban2017one], aims to predict dense masks for objects of novel classes using only a limited number of labeled support images. The mainstream prototype-based methods [dong2018few, li2021adaptive, wang2019panet] perform segmentation by measuring the similarity between the query features and representative support prototypes incorporating various improvements [siam2020weakly, liu2020dynamic, zhang2021self, zhuge2021deep]. The affinity-based methods [lu2021simpler, zhang2021few, peng2023hierarchical, min2021hypercorrelation, tian2020prior] establish detailed dense correspondence between query and support features through feature concatenation and leverage a learnable CNN or transformer module for segmentation prediction. Recently, foundation models like SAM [kirillov2023segment] present a novel opportunity for few-shot segmentation [liu2023matcher, zhang2023personalize], due to their remarkable transfer capability on tasks and data distributions beyond the training scope. However, these methods do not consider the domain shifts problem, leading to poor generalization performance when encountering new domains during testing.
@@
Cross-Domain Few-Shot Semantic Segmentation. Cross-domain few-shot semantic segmentation has recently received increasing attention. PATNet [lei2022cross] introduces a feature transformation layer that seamlessly maps query and support features across diverse domains into a unified feature space, effectively tackling the intra-domain knowledge preservation issue in CD-FSS. RD [wang2022remember] utilizes a memory bank to reinstill meta-knowledge from the source domain, thereby improving generalization performance in the target domain. Subsequently, DARNet [fan2023darnet] and RestNet [huang2023restnet] approach the problem from distinct perspectives, focusing on bridging domain gaps through dynamic adaptation refinement and knowledge transfer, respectively. Inspired by these pioneering efforts, PMNET [chen2024pixel] presents a comprehensive solution capable of addressing both in-domain and cross-domain FSS tasks concurrently by capturing pixel relations within each support-query pair. Unlike previous CD-FSS approaches, our method effectively addresses domain shifts in CD-FSS during inference by adaptively identifying and gradually adapting informative model structures on few-shot annotated support samples, and is capable of seamlessly adapting current few-shot segmentation methods to address domain shifts problem.
@@
Test-Time Training. Normally, once a well-trained model is deployed, it remains static without further alterations. In contrast, test-time training (TTT) [sun2020ttt] adapts models to the deployment scenario by leveraging unlabeled data available at test time. The mainstream self-supervised learning-based methods [liu2021ttt++, wang2021tent, liang2020do, goyal2022test, gandelsman2022test] leverage the available unlabeled test data to facilitate model adaptation to the target domain using self-supervised learning techniques. The feature alignment-based methods [su2022revisiting, jung2023cafa, wang2023feature] attempt to rectify the feature representations for the target domain. Some works attempt to apply TTT to address the semantic segmentation problem. For instance, MM-TTA [shin2022mmtta] utilizes multiple modalities to provide reciprocal TTT self-supervision for 3D semantic segmentation. Similarly, CD-TTA [song2022cdtta] explores domain-specific TTT for urban scene segmentation using an online clustering algorithm. OCL [zhang2024test] proposes an output contrastive loss to stabilize the TTT adaptation process for extreme class imbalance and complex decision spaces in semantic segmentation. These methods typically adjust fixed hand-selected trainable structures on one single test image for different domains. In contrast, our method dynamically adapt domain-specific informative structures by learning domain characteristics from few-shot labeled support samples.
@@ 3 Method
Cross-Domain Few-Shot Segmentation (CD-FSS) aims to transfer knowledge learned from the source domain to new categories in unseen target domains using minimal annotated supports. The model is typically trained on the source domain and then evaluated on target domains, ensuring no label space overlap between the source and target domains.
@@ 3.1 Baseline Methods
Few-Shot Segmentation Model. Mainstream few-shot segmentation model can be formulated as follows: The input support and query images $\{I_{s},I_{q}\}$ are processed by a weight-shared backbone to extract features $\{\mathcal{F}_{s},\mathcal{F}_{q}\}$:
@@
where $f$ denotes the image encoder with parameters $\theta$. Then, the support features $\mathcal{F}_{s}$ and groundtruth masks $\mathcal{M}_{s}$ are fed into the masked average pooling layer (MAP) to generate support prototypes $\mathcal{P}_{s}$. The prediction is made by measuring the cosine similarity between $\mathcal{P}_{s}$ and $\mathcal{F}_{q}$.
@@
3:for $n$ from $1$ to K1K-1 do $\triangleright$ PSA with increasing support shots. See Eqn. 7
4:  Extract features $\mathcal{F}$ for all samples using the updated model $\mathit{f}$ $\triangleright$ See Eqn. 1
@@
9:  Compute average loss $\mathcal{L}_{\texttt{PSA},n}$ for support shots $n$ $\triangleright$ See Eqn. 5
@@
11:  Update model $\mathit{f}$: $\theta_{\texttt{tr},n-1}^{*}$ $\rightarrow$ $\theta_{\texttt{tr},n}^{*}$ $\triangleright$ See Eqn. 6
@@
where BCE is the binary cross entropy loss and $\mathcal{M}_{q}$ is the groundtruth mask of the pseudo query image. Eventually, we train the model by optimizing the loss: $\theta^{*}=\underset{\theta}{\arg\min}\mathcal{L}^{i}_{\texttt{T}}(\mathcal{P}_{s}^{i},\mathcal{F}_{q}^{i},\mathcal{M}_{q}^{i};\theta),$ where $\theta$ denotes the trainable parameters of the model and $\theta^{*}$ denotes the updated model parameters after training. To prevent overfitting, we follow the common practice [boudiaf2021few, he2020momentum] to train only the final convolutional layer of the model during test-time training.
@@ 3.2 Informative Structure Identification
Empirical Fisher Information. However, directly computing the Fisher Information Matrix for the pre-trained CD-FSS model involves significant computational overhead due to the $|\theta|\times|\theta|$ scale. Thus, we simplify FIM for CD-FSS, inspired by the parameter importance metric used in continual learning [kirkpatrick2017overcoming]. Specifically, we concentrate on the support samples $K$ in the target domain and utilize the diagonal elements of the “Empirical Fisher” to evaluate the importance of the pre-trained model parameters for cross-domain tasks. Specifically, for the $l$-th convolutional layer, we derive the empirical Fisher information of its $u$-th trainable parameters as $F_{\theta_{l,u}}=\frac{1}{|K|}\sum_{j=1}^{|K|}\left(\frac{\partial\log p_{\theta}(y_{j}|x_{j})}{\partial\theta_{l,u}}\right)^{2}$. Correspondingly, a relatively large value of $F_{\theta_{l,u}}$ indicates that the parameter $\theta_{l,u}$ is crucial for the cross-domain task.
@@
Model layers with higher structure Fisher scores are typically more important for model training [liu2021sparse] because of their greater contribution to the optimization process. Updating only the informative model structures preserves the model’s ability to fit few-shot data and regularizes training to mitigate the risk of overfitting. Therefore, we select the model layer with the highest structure Fisher score and update its parameters $\theta_{\texttt{tr}}$ for model structure adaptation during inference, while freezing all other parameters to minimize the risk of overfitting in few-shot scenarios:
$\theta_{\texttt{tr}}=\theta_{l^{*}},{\mathrm{where}}\ l^{*}=\underset{l}{\arg\max}\{F^{*}_{\theta_{l}}\}.$ (4)
@@ 3.3 Progressive Structure Adaptation
Table 1: Quantitative comparison results on the CD-FSS benchmark. The models are trained on the Pascal VOC source domain dataset and evaluated on four datasets exhibiting distinct domain shifts. The best results are highlighted with bold. The $\dagger$ means our reproduced results. The $\ddagger$ means using the ViT-base backbone. All the in-domain FSS methods are directly applied to CD-FSS without source domain retraining (✗), thus suffering from the domain shifts problem. All other CD-FSS methods requires source domain retraining (✓) their models on the source domain to enhance domain generalization. Our method can be directly integrated with the well-trained FSS method (e.g., SSP) without source domain retraining (✗).
@@
PGNet [zhang2019pyramid] ICCV 2019 10.7 12.4 21.9 21.3 34.0 23.0 62.4 62.7 32.2 31.1
PANet [wang2019panet] ICCV 2019 36.6 45.4 25.3 34.0 57.8 69.3 69.2 71.7 47.2 55.1
CaNet [zhang2019canet] CVPR 2019 22.3 23.1 25.2 28.2 28.4 28.6 70.7 72.0 36.6 38.0
RPMMs [yang2020prototype] ECCV 2020 13.0 13.5 18.0 20.0 30.1 30.8 65.1 67.1 31.6 32.9
PFENet [tian2020prior] TPAMI 2020 16.9 18.0 23.5 23.8 27.2 27.6 70.9 70.5 34.6 35.0
RePRI [boudiaf2021few] CVPR 2021 25.0 27.4 23.3 26.2 65.1 65.5 71.0 74.2 46.1 48.3
HSNet [min2021hypercorrelation] ICCV 2021 29.7 35.1 31.2 35.1 51.9 54.4 77.5 81.0 47.6 51.4
SSP [fan2022self] ECCV 2022 42.3 50.4 33.0 47.0 74.9 75.5 77.1 79.1 56.8 63.0
ABCDFSS [herzog2024adapt] CVPR 2024 42.6 49.0 45.7 53.3 79.8 81.4 74.6 76.2 60.7 65.0
PATNet [lei2022cross] ECCV 2022 37.9 43.0 41.2 53.6 66.6 70.2 78.6 81.2 56.1 62.0
PMNet [chen2024pixel] WACV 2024 37.1 41.6 51.2 54.5 70.4 74.0 84.6 86.3 60.8 64.1
APSeg [he2024apseg] CVPR 2024 35.9 40.0 45.4 54.0 84.1 84.5 79.7 81.9 61.3 65.1
DR-Adapter [su2024domain] CVPR 2024 41.3 50.1 40.8 48.9 82.4 82.3 79.1 80.4 60.9 65.4
DMTNet [chen2024cross] IJCAI 2024 40.1 51.2 43.6 52.3 73.7 77.3 81.5 83.3 59.7 66.0
@@
Hierarchical Training Sample Construction. We begin by enhancing the utilization of few-shot support data, cyclically designating each support image as a pseudo query image to generate multiple support-query training pairs $\{(S\setminus S^{i}_{q},S^{i}_{q})\}_{i=1}^{K}$. To conserve computational resources, we first extract features from all support images, then compute the losses for each constructed training pair based on these features, and finally perform back-propagation to update model parameters using the averaged loss $\frac{1}{K}\sum_{i=1}^{K}\mathcal{L}_{\texttt{T}}^{i}$ across all training pairs. Next, we vary the number of support shots from $1$ to K1K-1 to construct hierarchical training pairs. Specifically, for each support shot number $n$ ($n\leq K-1$), we construct training pairs containing $n$ support samples. Thus, we progressively increase the number of support samples to construct training pairs, optimizing the utilization of few-shot support data and enabling the model to gradually handle domain shifts during inference.
Progressive Structure Adaptation. To better address domain shifts in cross-domain tasks during test-time training, we propose a Progressive Structure Adaptation (PSA) module that trains the model using HTC-constructed hierarchical training pairs. The PSA method progressively trains the model with an increasing number of support shots from $1$ to K1K-1, gradually reducing the domain gap. Specifically, for the support shot number $n$, the training loss is:
@@
where $\theta_{n-1}^{*}$ and $\theta_{n}^{*}$ denote the updated model parameters trained with HTC-constructed training samples with n1n-1 and $n$ support shots, respectively. Specifically, when $n=1$, the $\theta_{n-1}^{*}$ represents the original parameters $\theta$ of the FSS model.
Consequently, we progressively update the model parameters by gradually increasing the number of support shots from $1$ to K1K-1, as follows: $\theta_{1}^{*}\rightarrow\theta_{2}^{*}\rightarrow\dots\rightarrow\theta_{n}^{*}\rightarrow\dots\rightarrow\theta_{K-1}^{*}.$ This approach optimizes the use of limited support data to progressively mitigate domain shifts during inference on the support set. Hence, we term it as Progressive Structure Adaptation (PSA) module.